A hollow sphere of inner radius r and outer radius R floats half submerged in a liquid with density (P). If the density of the material of which the sphere is made is Pi, find an expression of the inner radius r in terms of R, Pi and P
1
Expert's answer
2020-06-29T14:06:19-0400

The volume submerged under water (water density is "\rho") is half the total volume of the sphere:


"V=\frac{V_R}{2}=\frac{2}{3}\pi R^3."

The buoyancy force for this condition is


"F_b=\rho gV=\frac{2}{3}\pi R^3\rho g."

The force of gravity acting on the sphere consists of the force of gravity of the sphere's material (material density is "\rho_i") assuming that there is vacuum inside:


"F_g=mg=\rho_iV_mg,"

where "V_m" - volume of the sphere's material:


"V_m=V_R-V_r=\frac{4}{3}\pi(R^3-r^3)."

Thus:


"F_g=\frac{4}{3}\pi(R^3-r^3)\rho_i g."

Since the sphere is floating, "F_b=F_g":


"\frac{2}{3}\pi R^3\rho g=\frac{4}{3}\pi(R^3-r^3)\rho_i g,\\\space\\
r=R\sqrt[3]{1-\frac{\rho}{2\rho_i}}."